# Derailer-Rerailer: Adaptive Verification for Efficient and Reliable Language Model Reasoning

### v3

@@ Abstract
Large Language Models (LLMs) have shown impressive reasoning capabilities, yet existing prompting methods face a critical trade-off: simple approaches often struggle with complex tasks and reasoning stability, while more sophisticated methods require multiple inferences and substantial computational resources, limiting their practical deployment. To address this challenge, we propose Derailer-Rerailer, a novel framework that adaptively balances reasoning accuracy and computational efficiency. At its core, our framework employs a lightweight Derailer mechanism to assess reasoning stability and selectively triggers an advanced Rerailer verification process only when necessary, thereby optimizing computational resource usage. Extensive evaluation across both open and closed-source models on more than 20 categories of mathematical, symbolic, and commonsense reasoning tasks demonstrates our framework’s effectiveness: Derailer-Rerailer achieves significant accuracy improvements (8-11% across various reasoning tasks) while maintaining 2-3 times better efficiency than existing verification methods, with particularly strong performance in mathematical and symbolic reasoning, offering a practical solution for enhancing LLM reasoning reliability while significantly reducing computational overhead  [^1] Code available at https://anonymous.4open.science/r/rerailer-0031/Readme.md.
@@ 1 Introduction
These complex prompting methods, however, face a critical trade-off between accuracy and computational efficiency. Such approaches dramatically increase computational overhead—For example in Self Consistency Wang et al. (2023) a single problem requires 40 model calls and results in thousands of tokens Li et al. (2024). Although this proves to be necessary in some complex situations, a limitation lies in their uniform application, indiscriminately deploying these expensive procedures across all queries regardless of whether the problem requires such intensive operations. This computational burden becomes particularly problematic in LLM applications where inference time is as crucial Snell et al. (2024) as accuracy, such as real-time clinical supports which requires both reliable reasoning and low latency Umerenkov et al. (2023), highlighting the need for an adaptive prompting method that can selectively apply computational resources to queries needed.
@@ 2.1 Emulating Human Cognitive Systems in Language Model Prompting
Iterative Prompting: Another category is iterative prompting which embraces a more deliberate approach. It either generates parallel independent samples (like Self-Consistency Wang et al. (2023) and use some sort of aggeration to obtain the final answer, or sequentially evaluates and refines the reasoning path based on the reward function (like chain of verification Dhuliawala et al. (2023):
@@
where $V$ represents reward functions that leverage additional model calls to refine the initial solutions or explore other indepedent solution paths. With a better reasoning performance, the computational cost of iterative prompting increases significantly scaled by both the number of attempts $k$ and the cost of verification $V$, which is usually done using LLM as a judge Gu et al. (2025).
The fact that human cognition does not necessitate deliberate processing for all decisions suggests a parallel for language models: computationally intensive refinement may not always be advantageous. This motivates the central question: Under what conditions is the investment in additional computation for deliberate reasoning warranted? Our proposed answer lies in analyzing the stability of the model’s reasoning process. We argue that the variability in LLM performance—ranging from high confidence to complete uncertainty, and including instances of knowledge without consistent application—suggests three distinct categories: Stable Success: When the model demonstrates consistent success with parallel sampling, usually in some single question like "what’s the result of 1 + 2", engaging in additional refinement thus becomes computationally inefficient. Stable Failure: When the model consistently fails regardless of the approach, indicating a fundamental knowledge or reasoning gap that additional computation cannot bridge. Unstable Reasoning: When the model shows potential but lacks consistency, producing a mix of successful and failed attempts. These cases represent ideal candidates for a system 2 style of thinking, where multiple attempts can help stabilize the reasoning process.
This stability-based framework motivates us to deisgn a principled approach for deploying computational resources efficiently, focusing intensive reasoning efforts on cases where the model demonstrates potential but requires additional guidance to achieve reliable performance.
@@ 2.2 Empirical Validation
As shown in Fig. 3, this analysis reveals two key patterns. First, regarding sampling efficiency, our experiments show that small samples already provide clear distinctions in consensus levels (the proportion of answers belonging to the majority class), with these patterns stabilizing at medium sampling sizes and larger samples offering diminishing returns. This aligns with recent findings (Wan et al., 2024a) on the effectiveness of reduced sampling in self-consistency approaches, which suggests a small set of samples determines the solvability of question. Second, we observe distinct peaks in both distributions: while questions with correct reasoning (orange) show high stability as expected, we also find a notable peak for cases of incorrect reasoning (blue) where questions consistently exceed the model’s capabilities. This suggests that sampling just a few examples can effectively identify a question’s stability category, thus motivating the design of an adaptive prompting method that adjusts computational investment based on reasoning stability, as suggested in Table 1.
@@ 3.1 Derailer: A Mechanism to Filter Stabilized Reasoning
The effectiveness of Derailer relies on two key conditions: First, the proportion of cases with stable reasoning should be substantial - a condition satisfied by modern LLMs which tend to be either consistently right or consistently wrong on many tasks as demonstrated in our preliminary studies and other work Wan et al. (2024a). Second, the stability check should be computationally efficient by keeping the number of samples $n$ low. By meeting these conditions, Derailer optimizes the trade-off between computational cost and answer quality, applying more intensive reasoning procedures only when the model’s unstable reasoning patterns suggest potential for improvement through iteration.
@@ 3.2 Rerailer: Stabilizing Inconsistent Reasoning
The core insight behind Rerailer is that unstable performance typically stems from error accumulation in multi-step reasoning, formalized as $P(y,S|x)=P(s_{1}|x)\prod_{i=2}^{n}P(s_{i}|s_{1:i-1},x)P(y|S,x)$, where errors at any intermediate step $s_{i}$ can propagate through conditional dependencies $P(s_{i}|s_{1:i-1},x)$, destabilizing the entire process even when the model possesses the fundamental capability. While approaches like Tree of Thoughts (ToT) similarly explore multiple decoding paths, Rerailer offers two key advantages: (1) the evaluation for each state employs pairwise comparisons between independent solutions rather than direct value assignment through few-shot learning, proving more reliable as models generally perform better at comparative judgments than absolute scoring; and (2) we employ a new search mechanism that dynamically explores branches only when comparisons yield ties, rather than maintaining a full tree exploration—enabling focused intervention precisely where reasoning instability occurs while preserving both accuracy and computational efficiency.
@@
Table 2: Comparison of reasoning methods across multiple models and datasets. All accuracy metrics are expressed in percentages with standard deviations (±) from multiple runs. Math Acc measures performance on mathematical reasoning problems. Symbolic Acc measures accuracy on symbolic reasoning tasks. Commonsense Acc reflects performance on common-sense reasoning tasks. Overall Acc is a combined measure across these categories. Tokens (K) indicates the total number of tokens used (in thousands). Acc/ K Token Gain (AGKT) shows how much the accuracy improves per 1K tokens over that model’s zero-shot CoT baseline. Values in parentheses next to the Overall Acc indicate the relative improvement compared to the zero-shot CoT baseline for each model.
@@
Claude-3.5-Sonnet Zero-shot CoT 68.3±1.8 77.6±1.2 72.2±1.5 72.7±1.5 0.108±0.03 -
Least-to-Most CoT 70.4±1.6 79.3±1.1 74.1±1.4 74.6±1.4 (+1.9) 0.372±0.04 5.11±0.6
Five-shot CoT 72.9±1.4 81.4±1.0 74.8±1.3 76.4±1.2 (+3.7) 0.518±0.08 7.14±0.8
Self-Consistency(SC) 77.8±0.8 86.5±0.6 75.2±0.7 79.8±0.7 (+7.1) 1.732±0.22 4.10±0.5
Derailer + SC 77.6±0.8 86.3±0.7 75.1±0.7 79.7±0.7 (+7.0) 0.762±0.08 9.19±0.7
Chain of Ver (CoVe) 78.2±1.0 86.8±0.8 75.4±0.9 80.1±0.9 (+7.4) 1.778±0.23 4.16±1.6
Derailer + CoVe 77.9±0.9 86.6±0.8 75.3±0.9 79.9±0.9 (+7.2) 0.793±0.11 9.08±0.8
Tree of Thought (ToT) 78.1±0.9 86.7±0.8 75.4±0.8 80.1±0.8 (+7.4) 2.245±0.25 3.30±0.6
Derailer + ToT 77.8±0.8 86.5±0.7 75.2±0.8 79.8±0.8 (+7.1) 0.845±0.12 8.40±0.7
Rerailer 78.0±0.8 86.7±0.7 75.3±0.7 80.0±0.7 (+7.3) 1.458±0.15 5.01±0.5
Derailer + Rerailer 80.8±0.5 89.7±0.4 75.9±0.5 82.1±0.5 (+9.4) 0.592±0.04 15.88±0.8
Llama-3.1 70B Zero-shot CoT 38.2±2.0 71.3±1.5 70.4±1.8 60.0±1.8 0.132±0.03 -
Least-to-Most CoT 39.6±1.8 72.4±1.4 72.3±1.6 61.4±1.6 (+1.4) 0.448±0.08 3.13±1.1
Five-shot CoT 41.8±1.5 74.7±1.2 72.9±1.4 63.1±1.4 (+3.1) 0.628±0.09 4.94±0.9
Self-Consistency(SC) 45.9±0.9 79.1±0.7 73.4±0.8 66.1±0.8 (+6.1) 1.932±0.17 3.16±1.0
Derailer + SC 45.7±0.9 78.8±0.8 73.3±0.8 65.9±0.8 (+5.9) 0.892±0.08 6.61±1.2
Chain of Ver (CoVe) 46.2±1.1 79.4±0.9 73.6±1.0 66.4±1.0 (+6.4) 1.972±0.28 3.25±1.7
Derailer + CoVe 46.0±1.0 79.2±0.9 73.5±0.9 66.2±0.9 (+6.2) 0.923±0.12 6.72±0.9
Tree of Thought (ToT) 46.1±1.0 79.3±0.8 73.5±0.9 66.3±0.9 (+6.3) 2.458±0.28 2.56±0.7
Derailer + ToT 45.9±0.9 79.1±0.8 73.4±0.8 66.1±0.8 (+6.1) 0.968±0.13 6.30±0.8
Rerailer 46.1±0.9 79.3±0.7 73.5±0.8 66.3±0.8 (+6.3) 1.652±0.14 3.81±0.8
Derailer + Rerailer 48.4±0.6 81.8±0.5 74.2±0.6 68.1±0.6 (+8.1) 0.648±0.10 12.50±1.1
GPT-4o-mini Zero-shot CoT 59.5±2.2 46.8±1.8 69.2±2.0 58.5±2.0 0.121±0.03 -
Least-to-Most CoT 62.4±2.1 49.1±1.7 71.8±1.9 61.1±1.9 (+2.6) 0.389±0.05 6.68±0.9
Five-shot CoT 63.9±1.8 50.8±1.5 72.4±1.7 62.4±1.7 (+3.9) 0.556±0.09 7.01±0.8
Self-Consistency (SC) 68.6±1.1 55.2±0.9 72.9±1.0 65.6±1.0 (+7.1) 1.795±0.14 3.96±0.9
Derailer + SC 68.3±1.1 54.9±1.0 72.8±1.0 65.3±1.0 (+6.8) 0.823±0.09 8.26±1.1
Chain of Ver (CoVe) 69.3±1.3 55.5±1.1 73.2±1.2 66.0±1.2 (+7.5) 1.842±0.25 4.07±1.8
Derailer + CoVe 68.9±1.2 55.3±1.0 73.1±1.1 65.8±1.1 (+7.3) 0.864±0.09 8.45±1.0
Tree of Thought (ToT) 69.2±1.2 55.4±1.0 73.1±1.1 65.9±1.1 (+7.4) 2.325±0.26 3.18±0.8
Derailer + ToT 68.8±1.1 55.2±0.9 73.0±1.0 65.7±1.0 (+7.2) 0.912±0.11 7.89±0.9
Rerailer 69.0±1.0 55.4±0.9 73.1±1.0 65.8±1.0 (+7.3) 1.539±0.17 4.74±0.9
Derailer + Rerailer 72.0±0.8 58.5±0.7 73.8±0.8 68.1±0.8 (+9.6) 0.639±0.09 15.02±1.2
@@ 5.1 Prompting LLMs for Reasoning
In-context Learning Dong et al. (2024), including few-shot Chain-of-Thought prompting Wei et al. (2022b) marked a breakthrough in eliciting reasoning capabilities from language models by encouraging them to articulate intermediate thought steps. Building upon this insight, researchers have developed various methods, including different thinking style decomposition Zhou et al. (2023), exploring multiple independent reasoning paths, Wang et al. (2023); Yao et al. (2024a), and sequential verification or refinement the generated rationale Madaan et al. (2023) Dhuliawala et al. (2023), or a mix of them Yao et al. (2024a). While other techniques which relies on external resource integration Shinn et al. (2023); Gao et al. (2023) and additional training can also enhance reasoning capabilities Trung et al. (2024) generally, our paper focuses on inference-time elicitation techniques for reasoning abilities already present in base models.
5.2 Inference Time Scaling and Efficiency
Despite strategic computation allocation during inference can enhance LLM reasoning Snell et al. (2024); Wan et al. (2024b), LLMs face significant efficiency challenges in elicitation Samsi et al. (2023); Wang et al. (2024a). .Current solutions rely on either carefully crafted prompts Wan et al. (2024a); Li et al. (2024) or additional training Setlur et al. (2024). In contrast, our Derailer approach directly leverages models’ ability to evaluate solution paths through answer solvability, enabling efficient test-time computation via adaptive refinement.

### v4

@@ Abstract
Large Language Models (LLMs) have shown impressive reasoning capabilities, yet existing prompting methods face a critical trade-off: simple approaches often struggle with complex tasks and reasoning stability, while more sophisticated methods require multiple inferences and substantial computational resources, limiting their practical deployment. To address this challenge, we propose Derailer-Rerailer, a novel framework that adaptively balances reasoning accuracy and computational efficiency. At its core, our framework employs a lightweight Derailer mechanism to assess reasoning stability and selectively triggers an advanced Rerailer verification process only when necessary, thereby optimizing computational resource usage. Extensive evaluation across both open and closed-source models on more than 20 categories of mathematical, symbolic, and commonsense reasoning tasks demonstrates our framework’s effectiveness: Derailer-Rerailer achieves significant accuracy improvements (8-11% across various reasoning tasks) while maintaining 2-3 times better efficiency than existing verification methods, with particularly strong performance in mathematical and symbolic reasoning, offering a practical solution for enhancing LLM reasoning reliability while significantly reducing computational overhead  [^1] Code available at https://github.com/wan19990901/CoT_rerailer.
@@ 1 Introduction
These complex prompting methods, however, face a critical trade-off between accuracy and computational efficiency. Such approaches dramatically increase computational overhead. For example, in Self-Consistency Wang et al. (2023), a single problem requires 40 model calls and results in thousands of tokens Li et al. (2024). While this computational intensity proves necessary for some complex situations, a key limitation lies in their uniform application—indiscriminately deploying these expensive procedures across all queries regardless of whether the problem actually requires such intensive operations. This computational burden becomes particularly problematic in real-world agentic applications Huang et al. (2024) that often involve various complex prompting methods, where inference time is as crucial Snell et al. (2024a) as accuracy (e.g. real-time clinical support agentic systems, which require both reliable reasoning and low latency Umerenkov et al. (2023); Wu et al. (2025a)). These constraints highlight the need for an adaptive prompting method that can selectively apply computational resources where they are most needed.
@@ 2.1 Emulating Human Cognitive Systems in Language Model Prompting
Iterative Prompting: Another category is iterative prompting which embraces a more deliberate approach. It either generates parallel independent samples (like Self-Consistency Wang et al. (2023) or Best of N sampling OpenAI (2022)) and use some sort of aggregation Wan et al. (2024) or selection Gu et al. (2025) mechanism to obtain the final answer, or sequentially evaluates and refines the reasoning path based on some evaluators (like Chain of Verification Dhuliawala et al. (2023) or Self-Refine Madaan et al. (2023)) Fomrally, this can be defined as:
@@
where $V$ represents reward functions that leverage additional model calls to refine the initial solutions or explore other indepedent solution paths. With a better reasoning performance, the computational cost of iterative prompting increases significantly scaled by both the number of attempts $k$ and the cost of verification $V$.
The fact that human cognition does not necessitate deliberate processing for all decisions suggests a parallel for language models: computationally intensive iterative prompting may not always be advantageous. This motivates the central question: Under what conditions is the investment in additional computation for deliberate reasoning (Iterated Prompting) warranted?
Our proposed answer lies in analyzing the stability of the model’s reasoning process. We argue that the variability in LLM performance—ranging from high confidence to complete uncertainty, and including instances of knowledge without consistent application—suggests three distinct categories: Stable Success: When the model demonstrates consistent success with parallel sampling, usually in some single question like "what’s the result of 1 + 2", engaging in additional refinement thus becomes computationally inefficient. Stable Failure: When the model consistently fails regardless of the approach, indicating a fundamental knowledge or reasoning gap that additional computation cannot bridge. Unstable Reasoning: When the model shows potential but lacks consistency, producing a mix of successful and failed attempts. These cases represent ideal candidates for a system 2 style of thinking, where multiple attempts can help stabilize the reasoning process.
@@ 2.2 Empirical Validation
As shown in Fig. 3, this analysis reveals two key patterns. First, regarding sampling efficiency, our experiments show that small samples already provide clear distinctions in consensus levels (the proportion of answers belonging to the majority class), with these patterns stabilizing at medium sampling sizes and larger samples offering diminishing returns. This aligns with recent findings (Wan et al., 2024) on the effectiveness of reduced sampling in self-consistency approaches, which suggests a small set of samples determines the solvability of question. Second, we observe distinct peaks in both distributions: while questions with correct reasoning (orange) show high stability as expected, we also find a notable peak for cases of incorrect reasoning (blue) where questions consistently exceed the model’s capabilities. This suggests that sampling just a few examples can effectively identify a question’s stability category, thus motivating the design of an adaptive prompting method that adjusts computational investment based on reasoning stability, as suggested in Table 1.
@@ 3.1 Derailer: A Mechanism to Filter Stabilized Reasoning
The effectiveness of Derailer relies on two key conditions: First, the proportion of cases with stable reasoning should be substantial - a condition satisfied by modern LLMs which tend to be either consistently right or consistently wrong on many tasks as demonstrated in our preliminary studies and other work Wan et al. (2024). Second, the stability check should be computationally efficient by keeping the number of samples $n$ low. By meeting these conditions, Derailer optimizes the trade-off between computational cost and answer quality, applying more intensive reasoning procedures only when the model’s unstable reasoning patterns suggest potential for improvement through iteration.
@@ 3.2 Rerailer: Stabilizing Inconsistent Reasoning
The core insight behind Rerailer is that unstable performance typically stems from error accumulation in multi-step reasoning, formalized as
$P(y,S|x)=P(s_{1}|x)\prod_{i=2}^{n}P(s_{i}|s_{1:i-1},x)P(y|S,x)$ (1)
where errors at any intermediate step $s_{i}$ can propagate through conditional dependencies $P(s_{i}|s_{1:i-1},x)$, destabilizing the entire process even when the model possesses the fundamental capability. While approaches like Tree of Thoughts (ToT) similarly explore multiple decoding paths, Rerailer offers two key advantages: (1) the evaluation for each state employs pairwise comparisons between independent solutions rather than direct value assignment through few-shot learning, proving more reliable as models generally perform better at comparative judgments than absolute scoring; and (2) we employ a new search mechanism that dynamically explores branches only when comparisons yield ties, rather than maintaining a full tree exploration—enabling focused intervention precisely where reasoning instability occurs while preserving both accuracy and computational efficiency.
@@
Table 2: Comparison of reasoning methods across multiple models and datasets. All accuracy metrics are expressed in percentages. Math Acc measures performance on mathematical reasoning problems. Symbolic Acc measures accuracy on symbolic reasoning tasks. Commonsense Acc reflects performance on common-sense reasoning tasks. Overall Acc is a combined measure across these categories. Tokens (K) indicates the total number of tokens used (in thousands). Acc/ K Token Gain (AGKT) shows how much the accuracy improves per 1K tokens over that model’s zero-shot CoT baseline. Values in parentheses next to the Overall Acc indicate the relative improvement compared to the zero-shot CoT baseline for each model.
@@
Claude-3.5-Sonnet Zero-shot CoT 68.3 77.6 72.2 72.7 0.108 -
Least-to-Most CoT 70.4 79.3 74.1 74.6 (+1.9) 0.372 5.11
Five-shot CoT 72.9 81.4 74.8 76.4 (+3.7) 0.518 7.14
Self-Consistency(SC) 77.8 86.5 75.2 79.8 (+7.1) 1.732 4.10
Derailer + SC 77.6 86.3 75.1 79.7 (+7.0) 0.762 9.19
Chain of Ver (CoVe) 78.2 86.8 75.4 80.1 (+7.4) 1.778 4.16
Derailer + CoVe 77.9 86.6 75.3 79.9 (+7.2) 0.793 9.08
Tree of Thought (ToT) 78.1 86.7 75.4 80.1 (+7.4) 2.245 3.30
Derailer + ToT 77.8 86.5 75.2 79.8 (+7.1) 0.845 8.40
Rerailer 78.0 86.7 75.3 80.0 (+7.3) 1.458 5.01
Derailer + Rerailer 80.8 89.7 75.9 82.1 (+9.4) 0.592 15.88
Llama-3.1 70B Zero-shot CoT 38.2 71.3 70.4 60.0 0.132 -
Least-to-Most CoT 39.6 72.4 72.3 61.4 (+1.4) 0.448 3.13
Five-shot CoT 41.8 74.7 72.9 63.1 (+3.1) 0.628 4.94
Self-Consistency(SC) 45.9 79.1 73.4 66.1 (+6.1) 1.932 3.16
Derailer + SC 45.7 78.8 73.3 65.9 (+5.9) 0.892 6.61
Chain of Ver (CoVe) 46.2 79.4 73.6 66.4 (+6.4) 1.972 3.25
Derailer + CoVe 46.0 79.2 73.5 66.2 (+6.2) 0.923 6.72
Tree of Thought (ToT) 46.1 79.3 73.5 66.3 (+6.3) 2.458 2.56
Derailer + ToT 45.9 79.1 73.4 66.1 (+6.1) 0.968 6.30
Rerailer 46.1 79.3 73.5 66.3 (+6.3) 1.652 3.81
Derailer + Rerailer 48.4 81.8 74.2 68.1 (+8.1) 0.648 12.50
GPT-4o-mini Zero-shot CoT 59.5 46.8 69.2 58.5 0.121 -
Least-to-Most CoT 62.4 49.1 71.8 61.1 (+2.6) 0.389 6.68
Five-shot CoT 63.9 50.8 72.4 62.4 (+3.9) 0.556 7.01
Self-Consistency (SC) 68.6 55.2 72.9 65.6 (+7.1) 1.795 3.96
Derailer + SC 68.3 54.9 72.8 65.3 (+6.8) 0.823 8.26
Chain of Ver (CoVe) 69.3 55.5 73.2 66.0 (+7.5) 1.842 4.07
Derailer + CoVe 68.9 55.3 73.1 65.8 (+7.3) 0.864 8.45
Tree of Thought (ToT) 69.2 55.4 73.1 65.9 (+7.4) 2.325 3.18
Derailer + ToT 68.8 55.2 73.0 65.7 (+7.2) 0.912 7.89
Rerailer 69.0 55.4 73.1 65.8 (+7.3) 1.539 4.74
Derailer + Rerailer 72.0 58.5 73.8 68.1 (+9.6) 0.639 15.02
@@ 5.1 Prompting LLMs for Reasoning
In-context Learning Dong et al. (2024), including few-shot Chain-of-Thought prompting Wei et al. (2022b), marked a breakthrough in eliciting reasoning capabilities from language models by encouraging them to articulate intermediate thought steps. Building upon this insight, researchers have developed various sophisticated methods to enhance reasoning performance. These approaches include different thinking style decompositions Zhou et al. (2023), exploring multiple parallel reasoning paths Wang et al. (2023); Yao et al. (2024a), and sequential verification or refinement of generated rationales Madaan et al. (2023); Dhuliawala et al. (2023), or combinations thereof Yao et al. (2024a); Snell et al. (2024a). Other techniques, particularly those related to LLM agents Wang et al. (2024a), rely on external knowledge or tool integration Shinn et al. (2023); Gao et al. (2023); Yao et al. (2023), while additional training methods can also enhance reasoning capabilities Trung et al. (2024). However, our paper focuses specifically on inference-time elicitation techniques for reasoning abilities already present in base models, which serve as the foundations for these more complex methods.
5.2 Efficiency in Large Language Model Inference
Despite significant advancements in reasoning capabilities through various prompting techniques, a critical challenge limiting the practical deployment of LLMs is their computational inefficiency Wang et al. (2025a). These models frequently exhibit unproductive behaviors, including generating redundant verifications after reaching correct answers or pursuing "fake thinking" patterns that simulate progress without meaningful problem-solving Wu et al. (2025b); Wang et al. (2025b), misallocating computational resources by applying uniform algorithms regardless of problem complexity or by inadequately distributing computation between simple and complex tasks Wan et al. (2024); Parashar et al. (2025); Yang et al. (2025). While existing solutions typically rely on sophisticated predictive strategies for budget allocation and algorithm selection Han et al. (2025); Snell et al. (2024b), we propose a simpler, more principled approach based on model’s consistency for the input queries with (1) a Derailer mechanism that enhances efficiency by terminating unproductive reasoning paths, and (2) a Rerailer component that recovers accuracy in cases where the Derailer incorrectly intervenes.